Write each of the following numerals in base 10. For base twelve, T and E represent the face values ten and eleven, respectively.
A.134five
B.11110two
C.83Etwelve

Found 2 solutions by Alan3354, greenestamps:
Answer by Alan3354(69443)

Write each of the following numerals in base 10. For base twelve, T and E represent the face values ten and eleven, respectively.
A.134five
If you mean base 5"
= 1*5^2 + 3*5^1 + 4
= 25 + 15 + 4
= 44
---------------------
B.11110two
If it's base 2:
= 30
Binary and Hex are useful.
Base 5 is not.
----------
C.83Etwelve
IDK what E is in base 12.

Answer by greenestamps(13367)



Here is a short algorithm for converting from another base to base 10:

(1) start with the first (leftmost) digit
(2) multiply by the base and add the next digit
(3) repeat step (2) until all digits have been used

A: 134 base 5 to base 10:

1*5 = 5; 5+3 = 8
8*5 = 40; 40+4 = 44

B: 11110 base 2 to base 5:

1*2 = 2; 2+1 = 3
3*2 = 6; 6+1 = 7
7*2 = 14; 14+1 = 15
15*2 = 30; 30+0 = 30

C: 83B base 12 to base 10 (in bases greater than 10, the digits after 9 are A, B, C, .... Use A and B instead of T and E for "digits" 10 and 11):

8*12 = 96; 96+3 = 99
99*12 = 1188; 1188+11 = 1199